# Remote Cooling of Spin-ensembles Coupled to a Spin-mechanical Hybrid Interface

### v1

@@ Appendix A Encoding a qubit into a macroscopic oscillator
The GKP code is a promising candidate for quantum error correction (QEC) in continuous variable systems [68], such as photonic systems and cavity modes. In addition, GKP states hold promise for enhanced sensing capabilities, allowing simultaneous detection of an oscillator’s displacements in both its position and momentum quadratures [64].
@@ A Oscillator displacements along the momentum and position quadratures
In the following, we assume that the spin ensemble is effectively decoupled from the rest of the system. The projectors given in Eq. (15) will then become just a displacement operator of the oscillator conditioning on the measurement result of the probe spin, as indicated by:
@@
The additional phase $\phi$ is introduced by rotating the ancilla spin by an angle $\phi$ before the measurement, as shown in the circuit in Fig. 6. The angle $\phi$ selects the effective measurement basis of the ancilla spin.
The proposed GKP encoding scheme involves repeated execution of this circuit, adaptively setting the $\phi$ angle based on previous measurement results. This is a phase estimation protocol that incrementally identifies the eigenvalue of $D(2\theta)$ while simultaneously projecting the oscillator into the corresponding eigenstate.
@@
For $\epsilon=0$, the filter function simplifies to $F(\epsilon,t)=2(n+2)$ so that $\theta=g_{0}t/\pi$. This leads to conditional displacements exactly along the position quadrature:
@@ B Experimental feasibility
In contrast to the oscillator cooling scheme developed in Ref. [58], which requires only a weak ancilla-oscillator coupling, the GKP encoding scheme is much more demanding as it requires a strong coupling strength. Current experimental configurations are clearly inadequate in this regard to implement the preparation scheme. Furthermore, the GKP encoding scheme has been extensively studied in Ref. [62].
Therefore, we choose not to perform detailed numerical simulations to evaluate the practicality of the GKP encoding scheme in this study. Instead, we briefly assess the challenges and discuss potential improvements to existing setups that could facilitate the preparation of GKP states in macroscopic mechanical oscillators. The time duration for each phase estimation round needs to be of the order of $1/g_{0}$, so that the displacement magnitude can be of the order of $|\theta|\sim O(\sqrt{\pi})$. Given the necessity of around 10 phase estimation cycles to yield an approximate GKP state of acceptable quality [62], the preparation scheme requires an overall time exceeding $\mathcal{O}(10/g_{0})$.
From an experimental point of view, it’s crucial that the coherence time of the oscillator is considerably long compared to the coupling strength $g_{0}$, ideally at least on the order of $100/g_{0}$. Diamond NMO with a quality factor $Q>10^{6}$ have been demonstrated at sub-Kelvin temperatures and mechanical damping rates of about a few kHz [69]. This implies that the oscillator coherence time needs to be significantly improved. Note that the coherence time of the NV electronic spin in cryogenic conditions can exceed 1 second through the application of dynamical decoupling [12]. This is far from sufficient, as the probe spin only needs to maintain coherence during each phase estimation round.

### v2

@@ Appendix C Encoding a qubit into a macroscopic oscillator
The GKP code is a promising candidate for quantum error correction (QEC) in continuous-variable systems [49], such as trapped-ion mechanical oscillators [56] and superconducting circuits [58]. In addition, GKP states hold the potential for enhanced sensing capabilities, allowing simultaneous detection of displacements in both the position and momentum quadratures of an oscillator [50].
@@ A Oscillator displacements along the momentum and position quadratures
The projectors given in Eq. (14) now become just a displacement operator of the oscillator, which depends on the measurement result of the probe spin, as indicated by:
@@
The additional phase $\phi$ is introduced by rotating the probe spin by an angle $\phi$ before the measurement, as shown in the circuit in Fig. 5. The proposed GKP encoding scheme involves repeated execution of this circuit, adaptively setting the angle $\phi$ based on previous measurement results. The angles $\phi$ select the effective measurement basis of the probe spin. This is a phase estimation protocol that incrementally identifies the eigenvalue of $D(2\theta)$ while simultaneously projecting the oscillator into the corresponding eigenstate.
@@
Our proposal is realized by carefully tuning the detuning $\epsilon$ and applying a certain number of inversion pulses to the probe spin. For $\epsilon=0$, the filter function simplifies to $F(\epsilon,t)=2(n+2)$ so that $\theta=g_{0}t/\pi$. This leads to conditional displacements exactly along the position quadrature:
@@ B Experimental feasibility
In contrast to the oscillator cooling scheme developed in Ref. [43], which requires only a weak probe-oscillator coupling, the GKP encoding scheme is much more demanding as it requires a strong coupling strength. In this respect, current experimental capabilities are clearly insufficient to implement the preparation scheme. Moreover, the GKP encoding scheme has been extensively studied in Ref. [48]. Therefore, we choose not to perform detailed numerical simulations to evaluate the practicality of the GKP encoding scheme in this study. Instead, we briefly assess the challenges and discuss potential improvements to existing setups that could facilitate the preparation of GKP states in macroscopic mechanical oscillators.
The time duration for each phase estimation round needs to be of the order of $1/g_{0}$, so that the displacement magnitude can be of the order of $|\theta|\sim O(\sqrt{\pi})$. Given the necessity of around 10 phase estimation cycles to yield an approximate GKP state of acceptable quality [48], the preparation scheme requires an overall time exceeding $\mathcal{O}(10/g_{0})$.
From an experimental point of view, it’s crucial that the coherence time of the oscillator is considerably long compared to the coupling strength $g_{0}$, ideally at least on the order of $100/g_{0}$. Diamond NMOs with a quality factor $Q>10^{6}$ have been demonstrated at sub-Kelvin temperatures and mechanical damping rates of a few kHz [59]. This implies that the coherence time of the vibrational mode must be significantly improved. Note that the coherence time of the NV electronic spin in cryogenic conditions can exceed 1 second by the application of dynamical decoupling [52]. This is far from sufficient, as the probe spin only needs to maintain coherence during each phase estimation round.